# Visual Explanations of Image-Text Representations via Multi-Modal Information Bottleneck Attribution

### v1

@@ 4.3 Localization Test
We quantitatively measure the effectiveness of our proposed attribution method by evaluating its accuracy in zero-shot detection for images. We binarize the saliency map such that the areas with scores higher than the threshold (75%) are assigned 1 while the rest are assigned 0. We denote the resulting binary map as $S_{\mathrm{pred}}$. We also construct a ground-truth binary map, $S_{\mathrm{gt}}$, using the bounding boxes provided by MS-CXR (Boecking et al., 2022), where the region inside the bounding boxes is assigned to 1 while the outside is assigned to 0. Note that some samples have multiple bounding boxes, and we consider all of them to test the method’s multi-occurrence detection ability. Then, we calculate the IoU (Intersection over Union) of $S_{\mathrm{pred}}$ and $S_{\mathrm{gt}}$. Namely, for images with a height of $n$ and a width of $m$, the score is calculated by
@@
We found that M2IB attribution attains an average IoU of 22.59% for this zero-shot detection task, outperforming all baseline models, as shown in Table 1. Recognizing that a localization score of 22.59% appears somewhat low in absolute terms, we briefly note that there are two potential causes that could lead to the low absolute values in the localization scores: (i) M2IB attribution indeed generates segmentation instead of bounding boxes, so evaluation by bounding boxes would underestimate the quality of the saliency map. (ii) The model under evaluation Endo et al. (2021) is not finetuned for detection and may only have learned a coarse-grained relationship between X-rays and medical notes.
@@
width= Methods GradCAM Selvaraju et al. (2016) Saliency Simonyan et al. (2014) KS Lundberg and Lee (2017) RISE Petsiuk et al. (2018) Chefer et al. (2021) Ours CC % Conf. Drop $\downarrow$ 4.96 $\pm$ 0.01 1.99 $\pm$ 0.01 1.94 $\pm$ 0.01 1.12 $\pm$ 0.01 1.63 $\pm$ 0.01 1.11 $\pm$ 0.01 image % Conf. Incr. $\uparrow$ 17.84 $\pm$ 0.08 22.95 $\pm$ 0.12 25.18 $\pm$ 0.28 35.72 $\pm$ 0.14 37.41 $\pm$ 0.12 41.55 $\pm$ 0.19 % ROAR+ $\uparrow$ 2.29 $\pm$ 0.41 6.88 $\pm$ 0.88 1.56 $\pm$ 0.88 3.15 $\pm$ 0.97 7.66 $\pm$ 0.55 10.59 $\pm$ 0.85 CC % Conf. Drop $\downarrow$ 2.19 $\pm$ 0.01 1.78 $\pm$ 0.01 1.71 $\pm$ 0.01 1.30 $\pm$ 0.01 1.06 $\pm$ 0.01 1.06 $\pm$ 0.01 text % Conf. Incr $\uparrow$ 29.71 $\pm$ 0.19 38.96 $\pm$ 0.15 46.87 $\pm$ 0.21 38.31 $\pm$ 0.48 38.42 $\pm$ 0.11 38.55 $\pm$ 0.20 % ROAR+ $\uparrow$ 43.23 $\pm$ 0.66 43.74 $\pm$ 0.65 47.46 $\pm$ 3.62 49.04 $\pm$ 1.12 53.57 $\pm$ 1.26 60.41 $\pm$ 1.12 % Conf. Drop $\downarrow$ 2.76 $\pm$ 0.03 0.81 $\pm$ 0.01 2.37 $\pm$ 0.04 3.94 $\pm$ 0.03 1.87 $\pm$ 0.02 0.55 $\pm$ 0.01 MSCXR % Conf. Incr. $\uparrow$ 12.64 $\pm$ 0.46 35.08 $\pm$ 0.44 10.24 $\pm$ 0.68 7.28 $\pm$ 0.44 21.44 $\pm$ 0.46 45.92 $\pm$ 0.70 image % ROAR+ $\uparrow$ 3.54 $\pm$ 0.80 25.46 $\pm$ 1.35 12.67 $\pm$ 1.02 16.79 $\pm$ 0.76 24.42 $\pm$ 1.19 38.7 $\pm$ 0.86 % Localization $\uparrow$ 5.56 $\pm$ 0.13 21.6 $\pm$ 0.16 7.77 $\pm$ 0.13 10.97 $\pm$ 0.24 21.65 $\pm$ 0.25 22.59 $\pm$ 0.14 MSCXR % Conf. Drop $\downarrow$ 2.26 $\pm$ 0.04 3.35 $\pm$ 0.03 2.4 $\pm$ 0.05 1.16 $\pm$ 0.02 2.93 $\pm$ 0.03 2.28 $\pm$ 0.04 text % Conf. Incr. $\uparrow$ 36.24 $\pm$ 0.54 18.88 $\pm$ 0.54 34.12 $\pm$ 0.77 57.2 $\pm$ 0.65 28.08 $\pm$ 0.34 35.48 $\pm$ 0.69 % ROAR+ $\uparrow$ 11.07 $\pm$ 0.62 15.79 $\pm$ 0.92 14.28 $\pm$ 1.09 12.09 $\pm$ 1.52 9.11 $\pm$ 0.6 16.31 $\pm$ 0.75
@@ 4.4 Degradation Tests
Drop in Confidence (Chattopadhay et al., 2018).    An ideal attribution method should only assign high scores to important features, thus we should not observe a drop in performance if only the high-attribution parts are allowed in the input. For images, we use point-wise multiplication of the saliency map and the image input. Since scaling token ids is meaningless, we use binarization similar to Wang et al. (2020) where only tokens with attribution scores in the top 50% are kept. We provide an example of distilled image and text in Figure 3. Formally, we define this score by
@@
Increase in Confidence (Chattopadhay et al., 2018).    Similarly, removing noisy information in the input might increase the model’s confidence. We compute
@@
Remove and Retrain + (ROAR+, an Extension of ROAR (Hooker et al., 2019)).   We finetune the base model on the degraded images and texts where the most important parts are replaced by uninformative values (i.e., channel means of images or padding tokens for texts, see Figure 3) and evaluated on a validation set of original inputs. [^3] The original ROAR also corrupts the validation data, which means that different methods might use varied training and validation datasets, making direct comparisons difficult. To ensure a fair comparison, we consistently use the original data as the validation set for all methods. If the attribution method is accurate, a sharp decrease in performance is expected because all useful features are removed, and the model cannot learn anything relevant from the degraded data. We split the testing dataset into 80% training data and 20% validation data. We use the same contrastive loss as used for pretraining CLIP and define the score by $(l_{c}-l_{o})/l_{o}$, where $l_{o}$ is the validation loss of retraining using the original data, and $l_{c}$ is the validation loss when retraining with corrupted data. We repeat the process five times and report the average score.
@@ 4.5 Sanity Check
We conduct a sanity check to ensure our method is, in fact, sensitive to model parameters. We follow the sanity check procedure proposed by Adebayo et al. (2018), where parameters in the model are randomized starting from the last to the first layer. As shown in Figure 4, M2IB passes the sanity check as the attribution scores of image pixels and text tokens change as the model weights change. Our method also produces more accurate saliency maps for finetuned models compared to pretrained models, which further confirms that the resulting attribution can successfully reflect the quality of the model. Since we insert the information bottleneck after a selected layer (layer 9 in this case), the randomization of this and previous layers appears to have a larger influence on the output.
Figure 4: Saliency Maps for Sanity Checks. “Finetuned” represents the model that is finetuned on MIMIC-CXR (Johnson et al., 2019), a Chest X-ray dataset. “Pretrained” represents pretrained CLIP (Radford et al., 2021) from OpenAI. “Projection” represents CLIP with randomized projection layer, which is the last layer of the image encoder or text encoder that projects image or text features into the shared embedding space. “Random” means that all parameters in the model are randomly initiated. The remaining columns represent models with weights randomized starting from the last to the given layer. The results suggest that the saliency maps of M2IB attribution are sensitive to model weights, as desired, meaning that M2IB passes the sanity check.
@@ 5 Discussion
Going beyond vision and language modalities, Girdhar et al. (2023) recently proposed a new multi-modal model, ImageBind, which aligns embeddings of five modalities to image embeddings through contrastive learning on pairs of images with each modality and uses a similar architecture as CLIP, where the encoder for each modality is a transformer. We applied M2IB attribution to pairs of six different modalities using ImageBind—which required minimal implementation steps—and provided qualitative examples in Appendix C to illustrate that M2IB attribution can be applied successfully to more than just vision and text modalities.
@@ Appendix A Experimental Details and Further Experimental Results
$\beta$ controls the relative importance of the compression term. The larger the $\beta$ is, the less information is allowed to flow through this layer. As shown in Figure 4(a), too large and too small $\beta$ generates similar attribution maps in terms of relative importance. However, too small $\beta$ allows nearly everything through the bottleneck, whereas too large $\beta$ nearly discards everything.
@@ Appendix B Attribution Maps for Additional Examples
Figure 6: Attribution maps for randomly picked examples from the MS-CXR chest X-ray dataset Boecking et al. [2022]. Note that the attribution score is assigned to each token, instead of each word, due to tokenization.
Figure 7: Attribution maps for randomly picked examples from the Conceptual Captions dataset Sharma et al. [2018].
Figure 8: Attribution maps for MRI and CT examples from the ROCO dataset Pelka et al. [2018]. The highlighted areas increase when $\beta$ decreases. Due to the lack of segmentation masks in ROCO, we are unable to perform a quantitative evaluation.
@@ Appendix C Extension to Additional Modalities With ImageBind
Our method can be easily extended to representation learning of modalities other than image and text, as long as features of different modalities are projected into a shared embedding space. We use ImageBind [Girdhar et al., 2023] as an example to showcase the effectiveness of our method in interpreting audio-image and audio-text representation learning. Since ImageBind aligns other modalities (audio, sensors that record depth (3D), thermal (infrared radiation), and inertial measurement units (IMU)) embedding to image embeddings through contrastive learning, the embeddings of all modalities are in one shared embedding space. Thus, we can insert the information bottleneck into certain layers of the feature encoders of ImageBind respectively and adopt the M2IB objective similar to Equation 6 and Equation 7.

### v2

@@ 4.3 Localization Test
We quantitatively measure the effectiveness of our proposed attribution method by evaluating its accuracy in zero-shot detection for images. We binarize the saliency map such that the areas with scores higher than the threshold (75%) are assigned 1 while the rest are assigned 0. We denote the resulting binary map as $S_{\mathrm{pred}}$. We also construct a ground-truth binary map, $S_{\mathrm{gt}}$, using the bounding boxes provided by MS-CXR [3], where the region inside the bounding boxes is assigned to 1 while the outside is assigned to 0. Note that some samples have multiple bounding boxes, and we consider all of them to test the method’s multi-occurrence detection ability. Then, we calculate the IoU (Intersection over Union) of $S_{\mathrm{pred}}$ and $S_{\mathrm{gt}}$. Namely, for images with a height of $n$ and a width of $m$, the score is calculated by
@@
We found that M2IB attribution attains an average IoU of 22.59% for this zero-shot detection task, outperforming all baseline models, as shown in Table 1. Recognizing that a localization score of 22.59% appears somewhat low in absolute terms, we briefly note that there are two potential causes that could lead to the low absolute values in the localization scores: (i) M2IB attribution indeed generates segmentation instead of bounding boxes, so evaluation by bounding boxes would underestimate the quality of the saliency map. (ii) The model under evaluation [8] is not finetuned for detection and may only have learned a coarse-grained relationship between X-rays and medical notes.
@@
width= Methods GradCAM [25] Saliency [28] KS [17] RISE [20] Chefer et al. [6] Ours CC % Conf. Drop $\downarrow$ 4.96 $\pm$ 0.01 1.99 $\pm$ 0.01 1.94 $\pm$ 0.01 1.12 $\pm$ 0.01 1.63 $\pm$ 0.01 1.11 $\pm$ 0.01 image % Conf. Incr. $\uparrow$ 17.84 $\pm$ 0.08 22.95 $\pm$ 0.12 25.18 $\pm$ 0.28 35.72 $\pm$ 0.14 37.41 $\pm$ 0.12 41.55 $\pm$ 0.19 % ROAR+ $\uparrow$ 2.29 $\pm$ 0.41 6.88 $\pm$ 0.88 1.56 $\pm$ 0.88 3.15 $\pm$ 0.97 7.66 $\pm$ 0.55 10.59 $\pm$ 0.85 CC % Conf. Drop $\downarrow$ 2.19 $\pm$ 0.01 1.78 $\pm$ 0.01 1.71 $\pm$ 0.01 1.30 $\pm$ 0.01 1.06 $\pm$ 0.01 1.06 $\pm$ 0.01 text % Conf. Incr $\uparrow$ 29.71 $\pm$ 0.19 38.96 $\pm$ 0.15 46.87 $\pm$ 0.21 38.31 $\pm$ 0.48 38.42 $\pm$ 0.11 38.55 $\pm$ 0.20 % ROAR+ $\uparrow$ 43.23 $\pm$ 0.66 43.74 $\pm$ 0.65 47.46 $\pm$ 3.62 49.04 $\pm$ 1.12 53.57 $\pm$ 1.26 60.41 $\pm$ 1.12 % Conf. Drop $\downarrow$ 2.76 $\pm$ 0.03 0.81 $\pm$ 0.01 2.37 $\pm$ 0.04 3.94 $\pm$ 0.03 1.87 $\pm$ 0.02 0.55 $\pm$ 0.01 MSCXR % Conf. Incr. $\uparrow$ 12.64 $\pm$ 0.46 35.08 $\pm$ 0.44 10.24 $\pm$ 0.68 7.28 $\pm$ 0.44 21.44 $\pm$ 0.46 45.92 $\pm$ 0.70 image % ROAR+ $\uparrow$ 3.54 $\pm$ 0.80 25.46 $\pm$ 1.35 12.67 $\pm$ 1.02 16.79 $\pm$ 0.76 24.42 $\pm$ 1.19 38.7 $\pm$ 0.86 % Localization $\uparrow$ 5.56 $\pm$ 0.13 21.6 $\pm$ 0.16 7.77 $\pm$ 0.13 10.97 $\pm$ 0.24 21.65 $\pm$ 0.25 22.59 $\pm$ 0.14 MSCXR % Conf. Drop $\downarrow$ 2.26 $\pm$ 0.04 3.35 $\pm$ 0.03 2.4 $\pm$ 0.05 1.16 $\pm$ 0.02 2.93 $\pm$ 0.03 2.28 $\pm$ 0.04 text % Conf. Incr. $\uparrow$ 36.24 $\pm$ 0.54 18.88 $\pm$ 0.54 34.12 $\pm$ 0.77 57.2 $\pm$ 0.65 28.08 $\pm$ 0.34 35.48 $\pm$ 0.69 % ROAR+ $\uparrow$ 11.07 $\pm$ 0.62 15.79 $\pm$ 0.92 14.28 $\pm$ 1.09 12.09 $\pm$ 1.52 9.11 $\pm$ 0.6 16.31 $\pm$ 0.75
@@ 4.4 Degradation Tests
Drop in Confidence [4].    An ideal attribution method should only assign high scores to important features, thus we should not observe a drop in performance if only the high-attribution parts are allowed in the input. For images, we use point-wise multiplication of the saliency map and the image input. Since scaling token ids is meaningless, we use binarization similar to [33] where only tokens with attribution scores in the top 50% are kept. We provide an example of distilled image and text in Figure 3. Formally, we define this score by
@@
Increase in Confidence [4].    Similarly, removing noisy information in the input might increase the model’s confidence. We compute
@@
Remove and Retrain + (ROAR+, an Extension of ROAR [11]).   We finetune the base model on the degraded images and texts where the most important parts are replaced by uninformative values (i.e., channel means of images or padding tokens for texts, see Figure 3) and evaluated on a validation set of original inputs. [^3] The original ROAR also corrupts the validation data, which means that different methods might use varied training and validation datasets, making direct comparisons difficult. To ensure a fair comparison, we consistently use the original data as the validation set for all methods. If the attribution method is accurate, a sharp decrease in performance is expected because all useful features are removed, and the model cannot learn anything relevant from the degraded data. We split the testing dataset into 80% training data and 20% validation data. We use the same contrastive loss as used for pretraining CLIP and define the score by $(l_{c}-l_{o})/l_{o}$, where $l_{o}$ is the validation loss of retraining using the original data, and $l_{c}$ is the validation loss when retraining with corrupted data. We repeat the process five times and report the average score.
@@ 4.5 Sanity Check
We conduct a sanity check to ensure our method is, in fact, sensitive to model parameters. We follow the sanity check procedure proposed by Adebayo et al. [1], where parameters in the model are randomized starting from the last to the first layer. As shown in Figure 4, M2IB passes the sanity check as the attribution scores of image pixels and text tokens change as the model weights change. Our method also produces more accurate saliency maps for finetuned models compared to pretrained models, which further confirms that the resulting attribution can successfully reflect the quality of the model. Since we insert the information bottleneck after a selected layer (layer 9 in this case), the randomization of this and previous layers appears to have a larger influence on the output.
Figure 4: Saliency Maps for Sanity Checks. “Finetuned” represents the model that is finetuned on MIMIC-CXR [13], a Chest X-ray dataset. “Pretrained” represents pretrained CLIP [21] from OpenAI. “Projection” represents CLIP with randomized projection layer, which is the last layer of the image encoder or text encoder that projects image or text features into the shared embedding space. “Random” means that all parameters in the model are randomly initiated. The remaining columns represent models with weights randomized starting from the last to the given layer. The results suggest that the saliency maps of M2IB attribution are sensitive to model weights, as desired, meaning that M2IB passes the sanity check.
@@ 5 Discussion
Going beyond vision and language modalities, Girdhar et al. [9] recently proposed a new multi-modal model, ImageBind, which aligns embeddings of five modalities to image embeddings through contrastive learning on pairs of images with each modality and uses a similar architecture as CLIP, where the encoder for each modality is a transformer. We applied M2IB attribution to pairs of six different modalities using ImageBind—which required minimal implementation steps—and provided qualitative examples in Appendix C to illustrate that M2IB attribution can be applied successfully to more than just vision and text modalities.
@@ Appendix A Experimental Details and Further Experimental Results
The hyperparameter $\beta$ controls the relative importance of the compression term. The larger the $\beta$ is, the less information is allowed to flow through this layer. As shown in Figure 5(a), too large and too small $\beta$ generates similar attribution maps in terms of relative importance. However, too small $\beta$ allows nearly everything through the bottleneck, whereas too large $\beta$ nearly discards everything.
@@ Appendix B Attribution Maps for Additional Examples
Figure 6: Attribution maps for randomly picked examples from the MS-CXR chest X-ray dataset [3]. Note that the attribution score is assigned to each token, instead of each word, due to tokenization.
Figure 7: Attribution maps for randomly picked examples from the Conceptual Captions dataset [26].
Figure 8: Attribution maps for MRI and CT examples from the ROCO dataset [19]. The highlighted areas increase when $\beta$ decreases. Due to the lack of segmentation masks in ROCO, we are unable to perform a quantitative evaluation.
@@ Appendix C Extension to Additional Modalities With ImageBind
Our method can be easily extended to representation learning of modalities other than image and text, as long as features of different modalities are projected into a shared embedding space. We use ImageBind [9] as an example to showcase the effectiveness of our method in interpreting audio-image and audio-text representation learning. Since ImageBind aligns other modalities (audio, sensors that record depth (3D), thermal (infrared radiation), and inertial measurement units (IMU)) embedding to image embeddings through contrastive learning, the embeddings of all modalities are in one shared embedding space. Thus, we can insert the information bottleneck into certain layers of the feature encoders of ImageBind respectively and adopt the M2IB objective similar to Equation 6 and Equation 7.